Six year ago,olu was five times as old as his son ope. How old is ope,if his age is now one-third of his father's present age.
Answer by Cromlix(4381)

Father - 6yrs = 5(Son - 6yrs)
F - 6 = 5(S - 6)
Son's age now = 1/3 of Father's
S = 1/3F
OR
F = 3S
Using:
F - 6 = 5(S - 6)
Replacing F with 3S
3S - 6 = 5(S - 6)
3S - 6 = 5S - 30
3S - 5S = -30 + 6
-2S = -24
2S = 24
S = 12
Ope is presently 12 years old.
Father's age is 36 years old.
6 years ago
Ope was 6 years old and his Father
was 5 times as old = 30 years old.
(5 times his son's age.)
Hope this helps.
:-)